OpenStudy (anonymous):

urn 1 contains 3 red balls and 4 black balls. urn 2 contains 4 red balls and 3 black balls. urn 3 contains 2 red balls and 6 black balls. if an urn is selected at random and a ball is drawn, find the probability it will be red

OpenStudy (anonymous):

Okay, so assuming we pick urn 1, what is the probability we get a red ball from it? And for urn 2? and urn 3? etc. Let's just start here.

OpenStudy (anonymous):

@carlo26 Can you do this?

OpenStudy (anonymous):

Remember that: \[ \Pr(A|B) = \frac{\Pr(A \cap B)}{\Pr(B)} \]

OpenStudy (anonymous):

1. 3/7 2. 4/7 3. 2/8

OpenStudy (anonymous):

Now, what is the probability that you urn 1 is selected?

OpenStudy (anonymous):

1/3 ?

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

By the way, what class are you taking? Do you understand what that formula I gave up there meant?

OpenStudy (anonymous):

intro to prob & stats.. no, i didn't lol

OpenStudy (anonymous):

so the answer is 1/3? lol

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

ok.. what do i do now, after i got that?

OpenStudy (anonymous):

Basically, you calculated the probability of picking a red ball from urn 1 IF you knew that a ball was picked from urn 1. If you multiply these probabilities, you will get the probability that you picked urn 1 and picked a red ball from it.

OpenStudy (anonymous):

So the probability of picking urn 1 AND picking a red ball is: \(\frac{3}{7} \times \frac{1}{3}\).

OpenStudy (anonymous):

Now you need the probability picking urn 2 AND picking a red ball... and the same for picking urn 3 AND picking a red ball

OpenStudy (anonymous):

Once you have calculated all three, the probability of picking a ball from any of them is adding those 3 together.

OpenStudy (anonymous):

i hate this stuff lol.. thanks for the great helping though!

OpenStudy (anonymous):

prob of urn 2 and picking a red ball is 1/3 and 4/7, right?

OpenStudy (anonymous):

multiplied

OpenStudy (anonymous):

yea.. so i just multiply all 3 of the things after i multiply them individually?

OpenStudy (anonymous):

no you add them up

OpenStudy (anonymous):

k

OpenStudy (anonymous):

thus reqd prob =1/3(3/7 +4/7 +2/8)=5/12